OpenStudy (anonymous):

Write the equation in standard form x^2+y^2-2x-6y+9=0

OpenStudy (cwrw238):

(x-1)^2 + (y-3)^2 + 19 = 0

OpenStudy (cwrw238):

is that what you require?

OpenStudy (cwrw238):

sorry - error its (x-1)^2 + (y-3)^2 - 1 - 9 + 9 = 0 so (x-1)^2 + (y-3)^2 = 1